# Graph Attention with Random Rewiring

### v1

@@ 3.4 Interpretations of GRASS
Noise injection has been widely known to be an effective regularizer [50]. For graph-structured data, it has been observed that randomly dropping edges has regularization effects [56]. By adding noise to the adjacency matrix that guides the propagation of information in message passing, GRASS is encouraged to be robust against structural noise, which is prevalent in graph-structured data [64].
@@
GRASS is a sparse Graph Transformer. Graph Transformers allow each node to aggregate information from other nodes through the attention mechanism, with a general definition [65] being
@@
Many GTs achieve sparsity by integrating [55] or generalizing [60] BigBird’s sparse dot-product attention. Meanwhile, GRASS achieves sparsity in a graph-native way: nodes that are unconnected in the rewired graph would naturally not attend to each other. Seeing GRASS as a Transformer, its attention mask would be $\mathbf{A}_{M}$ as defined in Equation 18. Moreover, there might be little computational benefit in using dot-product attention on graphs, as the attention matrix would have unstructured sparsity. Efficient implementations of dot-product attention require efficient multiplication of sparse matrices, which is difficult with the unstructured sparsity prevalent in graph-structured data [11].
@@ 4 Experiments
[…]
@@
Table 2: Performance on LRGB datasets. The performance of GRASS shown here is the mean $\pm$ s.d. of 8 runs. The best and second-best results are highlighted. Performance numbers other than that of GRASS are adapted from Gutteridge et al. [31], Ma et al. [46], and Shirzad et al. [60].
@@
Model Peptides-func Peptides-struct PascalVOC-SP AP $\uparrow$ MAE $\downarrow$ Macro F1 $\uparrow$ GCN 0.5930 $\pm$ 0.0023 0.3496 $\pm$ 0.0013 0.1268 $\pm$ 0.0060 GINE 0.5498 $\pm$ 0.0079 0.3547 $\pm$ 0.0045 0.1265 $\pm$ 0.0076 GatedGCN 0.5864 $\pm$ 0.0077 0.3420 $\pm$ 0.0013 0.2873 $\pm$ 0.0219 GatedGCN+RWSE 0.6069 $\pm$ 0.0035 0.3357 $\pm$ 0.0006 0.2860 $\pm$ 0.0085 DIGL+MPNN 0.6469 $\pm$ 0.0019 0.3173 $\pm$ 0.0007 0.2824 $\pm$ 0.0039 DIGL+MPNN+LapPE 0.6830 $\pm$ 0.0026 0.2616 $\pm$ 0.0018 0.2921 $\pm$ 0.0038 MixHop-GCN 0.6592 $\pm$ 0.0036 0.2921 $\pm$ 0.0023 0.2506 $\pm$ 0.0133 MixHop-GCN+LapPE 0.6843 $\pm$ 0.0049 0.2614 $\pm$ 0.0023 0.2218 $\pm$ 0.0174 DRew-GCN 0.6996 $\pm$ 0.0076 0.2781 $\pm$ 0.0028 0.1848 $\pm$ 0.0107 DRew-GCN+LapPE 0.7150 $\pm$ 0.0044 0.2536 $\pm$ 0.0015 0.1851 $\pm$ 0.0092 DRew-GIN 0.6940 $\pm$ 0.0074 0.2799 $\pm$ 0.0016 0.2719 $\pm$ 0.0043 DRew-GIN+LapPE 0.7126 $\pm$ 0.0045 0.2606 $\pm$ 0.0014 0.2692 $\pm$ 0.0059 DRew-GatedGCN 0.6733 $\pm$ 0.0094 0.2699 $\pm$ 0.0018 0.3214 $\pm$ 0.0021 DRew-GatedGCN+LapPE 0.6977 $\pm$ 0.0026 0.2539 $\pm$ 0.0007 0.3314 $\pm$ 0.0024 Transformer+LapPE 0.6326 $\pm$ 0.0126 0.2529 $\pm$ 0.0016 0.2694 $\pm$ 0.0098 SAN+LapPE 0.6384 $\pm$ 0.0121 0.2683 $\pm$ 0.0043 0.3230 $\pm$ 0.0039 GPS+LapPE 0.6535 $\pm$ 0.0041 0.2500 $\pm$ 0.0005 0.3748 $\pm$ 0.0109 Exphormer 0.6527 $\pm$ 0.0043 0.2481 $\pm$ 0.0007 0.3975 $\pm$ 0.0037 GRIT 0.6988 $\pm$ 0.0082 0.2460 $\pm$ 0.0012 - GRASS (ours) 0.6353 $\pm$ 0.0056 0.2447 $\pm$ 0.0009 0.5633 $\pm$ 0.0092
@@
It has been suggested that oversmoothing can be alleviated by having a shallow GNN, and oversquashing can be avoided by having wide hidden layers [62]. However, narrow-and-deep GNNs can have better expressivity than wide-and-shallow GNNs if the pitfalls of training deep GNNs, such as oversmoothing and oversquashing, can be avoided [42, 43]. This could explain the competitive performance of GRASS which we configured to be narrow and deep in our experiments, as specified in Appendix E. In addition, the datasets we tested are relatively small in sample size (<100k samples), which makes a good inductive bias more important [5, 46], highlighting the ability of GRASS to provide a strong inductive bias on graph-structured data while remaining computationally efficient.
@@
GRASS achieves its design goals with multiple techniques, whose importance is verified by the ablation studies shown in Table 3. In particular, the performance of GRASS with different values of $r$, the degree of the superimposed random regular graph, is compared in Table 4.
@@ Appendix A Pseudocode of the Permutation Model
12:     $A\leftarrow\textsc{RemoveSelfLoop}(A)$triangleright Remove self-loops from $A$
13:     $A\leftarrow\textsc{RemoveMultiEdge}(A)$triangleright Remove multi-edges from $A$
@@ Appendix C Ablation Study
Table 3: Ablation study results on different components of GRASS. This table shows the performance of each ablated model as the mean $\pm$ s.d. over 8 runs.
Model ZINC (MAE $\downarrow$) GRASS 0.0461 $\pm$ 0.0014       Remove RRWP encoding 0.0797 $\pm$ 0.0040       Remove DropKey 0.0549 $\pm$ 0.0125       Remove edge flipping 0.0477 $\pm$ 0.0010       PN-V $\rightarrow$ LN 0.0497 $\pm$ 0.0013       Sum pooling $\rightarrow$ Mean pooling 0.0514 $\pm$ 0.0018

### v2

@@ 3.4 Interpretations of GRASS
Noise injection is well known as an effective regularizer [51], and for graph-structured data, the random removal of edges has demonstrated regularization effects [57]. By adding noise to the adjacency matrix that guides the propagation of information in message passing, GRASS is encouraged to develop robustness against structural noise, which is prevalent in graph-structured data [66].
@@
GRASS is a sparse Graph Transformer. Graph Transformers allow each node to aggregate information from other nodes through graph attention mechanisms, with a general definition [67] being
@@
Many GTs achieve sparsity by integrating [56] or generalizing [61] BigBird’s sparse dot-product attention. Meanwhile, GRASS achieves sparsity in a graph-native way: attention is always local, so nonadjacent nodes in the rewired graph would naturally never attend to each other. Seeing GRASS as a Transformer, its attention mask would be $\mathbf{A}_{M}$ as defined in Equation 18. Moreover, there might be little computational benefit in using sparse dot-product attention on graphs-structured data, since its efficient implementation relies on efficient sparse matrix multiplication [76], which is challenging [28, 12] due to the lack of exploitable sparsity patterns in many real-world graphs.
@@ 4 Experiments
Model ZINC MNIST CIFAR10 PATTERN CLUSTER MAE $\downarrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ GCN 0.367 $\pm$ 0.011 90.705 $\pm$ 0.218 55.710 $\pm$ 0.381 71.892 $\pm$ 0.334 68.498 $\pm$ 0.976 GIN 0.526 $\pm$ 0.051 96.485 $\pm$ 0.252 $5$$\pm$ 1.527 85.387 $\pm$ 0.136 64.716 $\pm$ 1.553 GAT 0.384 $\pm$ 0.007 95.535 $\pm$ 0.205 64.223 $\pm$ 0.455 78.271 $\pm$ 0.186 70.587 $\pm$ 0.447 GatedGCN 0.282 $\pm$ 0.015 97.340 $\pm$ 0.143 67.312 $\pm$ 0.311 85.568 $\pm$ 0.088 73.840 $\pm$ 0.326 GatedGCN-LSPE 0.090 $\pm$ 0.001 - - - - PNA 0.188 $\pm$ 0.004 97.94 $\pm$ 0.12 70.35 $\pm$ 0.63 - - DGN 0.168 $\pm$ 0.003 - 72.838 $\pm$ 0.417 86.680 $\pm$ 0.034 - GSN 0.101 $\pm$ 0.010 - - - - CIN 0.079 $\pm$ 0.006 - - - - CRaW1 0.085 $\pm$ 0.004 97.944 $\pm$ 0.050 69.013 $\pm$ 0.259 - - GIN-AK+ 0.080 $\pm$ 0.001 - 72.19 $\pm$ 0.13 86.850 $\pm$ 0.057 - SAN 0.139 $\pm$ 0.006 - - - 76.691 $\pm$ 0.65 Graphormer 0.122 $\pm$ 0.006 - - - - K-Subgraph SAT 0.094 $\pm$ 0.008 - - 86.848 $\pm$ 0.037 77.856 $\pm$ 0.104 EGT 0.108 $\pm$ 0.009 98.173 $\pm$ 0.087 68.702 $\pm$ 0.409 86.821 $\pm$ 0.020 79.232 $\pm$ 0.348 Graphormer-URPE 0.086 $\pm$ 0.007 - - - - Graphormer-GD 0.081 $\pm$ 0.009 - - - - GPS 0.070 $\pm$ 0.004 - 72.298 $\pm$ 0.356 86.685 $\pm$ 0.059 78.016 $\pm$ 0.180 Exphormer - 98.55 $\pm$ 0.039 74.69 $\pm$ 0.125 86.74 $\pm$ 0.015 78.07 $\pm$ 0.037 GRIT 0.059 $\pm$ 0.002 98.108 $\pm$ 0.111 76.468 $\pm$ 0.881 87.196 $\pm$ 0.076 80.026 $\pm$ 0.277 GRASS (ours) 0.046 $\pm$ 0.001 98.905 $\pm$ 0.034 82.471 $\pm$ 0.319 90.997 $\pm$ 0.225 79.347 $\pm$ 0.087
@@
Table 2: Performance on LRGB datasets. The performance of GRASS shown here is the mean $\pm$ s.d. of 8 runs. The best and second-best results are highlighted. Performance numbers other than that of GRASS are adapted from Gutteridge et al. [32], Tönshoff et al. [63], Ma et al. [47], and Shirzad et al. [61]. *Tönshoff et al. [63] has improved the performance of these models via hyperparameter tuning.
@@
Model Peptides-func Peptides-struct PascalVOC-SP AP $\uparrow$ MAE $\downarrow$ Macro F1 $\uparrow$ GCN* 0.6860 $\pm$ 0.0050 0.2460 $\pm$ 0.0007 0.2078 $\pm$ 0.0031 GINE* 0.6621 $\pm$ 0.0067 0.2473 $\pm$ 0.0017 0.2718 $\pm$ 0.0054 GatedGCN* 0.6765 $\pm$ 0.0047 0.2477 $\pm$ 0.0009 0.3880 $\pm$ 0.0040 DIGL+MPNN 0.6469 $\pm$ 0.0019 0.3173 $\pm$ 0.0007 0.2824 $\pm$ 0.0039 DIGL+MPNN+LapPE 0.6830 $\pm$ 0.0026 0.2616 $\pm$ 0.0018 0.2921 $\pm$ 0.0038 MixHop-GCN 0.6592 $\pm$ 0.0036 0.2921 $\pm$ 0.0023 0.2506 $\pm$ 0.0133 MixHop-GCN+LapPE 0.6843 $\pm$ 0.0049 0.2614 $\pm$ 0.0023 0.2218 $\pm$ 0.0174 DRew-GCN 0.6996 $\pm$ 0.0076 0.2781 $\pm$ 0.0028 0.1848 $\pm$ 0.0107 DRew-GCN+LapPE 0.7150 $\pm$ 0.0044 0.2536 $\pm$ 0.0015 0.1851 $\pm$ 0.0092 DRew-GIN 0.6940 $\pm$ 0.0074 0.2799 $\pm$ 0.0016 0.2719 $\pm$ 0.0043 DRew-GIN+LapPE 0.7126 $\pm$ 0.0045 0.2606 $\pm$ 0.0014 0.2692 $\pm$ 0.0059 DRew-GatedGCN 0.6733 $\pm$ 0.0094 0.2699 $\pm$ 0.0018 0.3214 $\pm$ 0.0021 DRew-GatedGCN+LapPE 0.6977 $\pm$ 0.0026 0.2539 $\pm$ 0.0007 0.3314 $\pm$ 0.0024 Transformer+LapPE 0.6326 $\pm$ 0.0126 0.2529 $\pm$ 0.0016 0.2694 $\pm$ 0.0098 SAN+LapPE 0.6384 $\pm$ 0.0121 0.2683 $\pm$ 0.0043 0.3230 $\pm$ 0.0039 GPS+LapPE* 0.6534 $\pm$ 0.0091 0.2509 $\pm$ 0.0014 0.4440 $\pm$ 0.0065 Exphormer 0.6527 $\pm$ 0.0043 0.2481 $\pm$ 0.0007 0.3975 $\pm$ 0.0037 GRIT 0.6988 $\pm$ 0.0082 0.2460 $\pm$ 0.0012 - GRASS (ours) 0.6353 $\pm$ 0.0056 0.2447 $\pm$ 0.0009 0.5633 $\pm$ 0.0092
@@
It has been suggested that oversmoothing can be alleviated by having a shallow GNN, and oversquashing can be avoided by having wide hidden layers [64]. However, narrow-and-deep GNNs can have better expressivity than wide-and-shallow GNNs if the pitfalls of training deep GNNs, such as oversmoothing and oversquashing, can be avoided [43, 44]. This could potentially explain the competitive performance of GRASS which we configured to be narrow and deep in our experiments, as specified in Appendix E. In addition, the datasets we tested are relatively small in sample size (<100k samples), which makes a good inductive bias more important [6, 47], highlighting the ability of GRASS to provide a strong inductive bias on graph-structured data while remaining computationally efficient.
@@
GRASS achieves its design goals with multiple components, whose importance is evaluated in the ablation studies shown in Table 3. In particular, the performance of GRASS with different values of $r$, the degree of the superimposed random regular graph, is compared in Table 4.
@@ Appendix A Pseudocode of the Permutation Model
12:     $A\leftarrow\textsc{RemoveSelfLoop}(A)$triangleright Remove self-loops from the adjacency list
13:     $A\leftarrow\textsc{RemoveMultiEdge}(A)$triangleright Remove multi-edges from the adjacency list
@@ Appendix C Ablation Study
Table 3: Ablation study results on different components of GRASS. This table shows the performance of each ablated model as the mean $\pm$ s.d. over 8 runs. *The absolute difference between the LapPEs (8 eigenvalues) of neighboring nodes are used as edge encodings, as proposed by Kreuzer et al. [40].
Model ZINC MAE $\downarrow$ GRASS 0.0461 $\pm$ 0.0014       Remove RRWP encoding 0.0797 $\pm$ 0.0040       RRWP encoding $\rightarrow$ relative LapPE* 0.0708 $\pm$ 0.0032       Remove DropKey 0.0549 $\pm$ 0.0125       Remove edge flipping 0.0477 $\pm$ 0.0010       PN-V $\rightarrow$ LN [3] 0.0497 $\pm$ 0.0013       Sum pooling $\rightarrow$ Mean pooling 0.0514 $\pm$ 0.0018